# Mohana Prasad Sathya Moorthy, Jeff Nichols, Yinfei Yang, Zhe Gan Apple

### v1

@@ 1 Introduction
One prominent effort in this space is Ferret-UI (You et al., 2024), which has advanced the field of referring and grounding UIs. However, though taking an any-resolution approach (Liu et al., 2024a), Ferret-UI is constrained by a fixed grounding resolution (i.e., 336$\times$672 and 672$\times$336), and focuses on single-type platforms (i.e., mobile devices including iPhone and Android), limiting its applicability in the context of today’s highly diverse platform landscape. For example, as illustrated in Figure 1, one notable difference among these four exemplified platforms is resolution, the native resolution of iPhone differs from that of iPad, Web UI and also AppleTV, directly applying Ferret-UI across these platforms presents significant challenges. Another major challenge is the lack of platform-specific, high-quality data, given different platforms. Though Ferret-UI’s approach to training data generation can be extended to these platforms, it primarily relies on text-based GPT-4 prompting, where bounding boxes are represented in a purely textual format. This absence of visual input and spatial relationships between UI elements diminishes the quality of training data, which in turn limits the performance and effectiveness of the resulting model.
Figure 1: Real examples of a single Ferret-UI 2 model interacting with four different platforms (iPhone, iPad, Webpage, and AppleTV) for UI understanding. Refer to Appendix D for more examples including multi-step interactions.
@@
Second, Ferret-UI 2 supports high-resolution image encoding via the any-resolution method (Liu et al., 2024a; Zhang et al., 2024c). However, going beyond that, we introduce an enhanced adaptive gridding approach to maintain perception capabilities at the original resolution of the UI screenshot, ensuring more accurate recognition of visual elements. By leveraging human-collected bounding box annotations, we enhance referring and grounding precision, which allows for more detailed understanding of UI components and their relationships.
@@ 2 Related Work
Multi-Platform UI Agents. Multi-platform UI agents have emerged to address the growing complexity of device ecosystems, supporting a variety of devices and platforms, including mobile, web, and desktop environments. Zheng et al. (2024a); Cheng et al. (2024) demonstrated GPT-4V as a generalist agent when grounded. Recent works like OmniACT (Kapoor et al., 2024) support both desktop and web interfaces, while CogAgent (Hong et al., 2023) supports UI navigation on both PC webpages and Android devices. Furthermore, Mind2Web (Deng et al., 2024) and Mobile-Agent (Wang et al., 2024b) are also notable research works that enable agents to operate seamlessly across different platforms. Mobile-Agent V2 (Wang et al., 2024a) features Harmony OS and Android OS for non-English and English scenarios. Ferret-UI (You et al., 2024) focuses on mobile UI understanding for both Android and iPhone screenshots using multimodal LLMs (McKinzie et al., 2024; Zhang et al., 2024b), with a focus on referring and grounding capabilities. These agents aim to perform more complex, user-intent-based interactions across multiple device types, paving the way for truly generalist multimodal agents.
UI-Agent Benchmarks. The evaluation of UI agents requires specialized benchmarks to test various aspects of UI interaction, including task execution, navigation, and interaction understanding. Rico (Deka et al., 2017) remains a foundational dataset for mobile app interaction, while Mobile-Env (Zhang et al., 2023), AndroidEnv (Toyama et al., 2021), AndroidWorld (Rawles et al., 2024a), Android in-the-Wild (Rawles et al., 2024b), AndroidControl (Li et al., 2024b), and AMEX (Chai et al., 2024) provide benchmarks for mobile device control. Windows Agent Arena (Bonatti et al., 2024) introduces benchmark focusing on PC windows environment. OSWorld (Xie et al., 2024) takes a broader approach by providing benchmarks for agents in real computer environments, including Ubuntu, MacOS, and Windows. Web-based interaction benchmarks include WebSRC (Chen et al., 2021), Mind2Web (Deng et al., 2024) and WebCanvas (Pan et al., 2024) focusing on structural reading comprehension and task execution in web environments. More recently, MobileAgentBench (Wang et al., 2024c) and VisualWebBench (Liu et al., 2024b) have introduced taxonomies designed to evaluate the performance of multimodal agents across both mobile and web interfaces. VisualAgentBench (Liu et al., 2024c) expands this with a focus on multimodal LLMs as visual foundation agents. GUI Odyssey (Lu et al., 2024) provides benchmarks for cross-App navigation. GUI-World (Chen et al., 2024a) pioneers in covering multi-platform benchmarking for GUI agents, while CRAB (Xu et al., 2024) further tests cross-environment tasks for GUI agents. These benchmarks contribute to a growing need for unified, multi-platform evaluations that can assess UI agents’ adaptability, precision, and efficiency.
@@ 3.1 Dataset Construction
  • iPhone, iPad, and AppleTV: We use human-collected iPhone, iPad and AppleTV data under diverse usage scenarios and human-annotated widget bounding box coordinates and labels. To save annotation costs, we do not collect text annotations; instead, text bounding boxes are replaced by screen-wide OCR-detected text and bounding boxes using an OCR confidence threshold of 0.5.
@@
  • Android: The Android data for screenshots, bounding boxes and text annotations is transformed from the RICO dataset (Deka et al., 2017). Similar to the WebUI dataset, we also perform picture-only OCR to complete the missing text annotations in picture widgets.
@@
Table 1: A summary of datasets across various platforms used to train Ferret-UI 2. The screenshot resolution statistics are shown in Appendix E.
@@
Figure 3: Example of set-of-mark visual prompting (Yang et al., 2023) (left) and one of its generated advanced task training example (right).
Despite the different types of label spaces from various sources, we filter out bounding boxes associated with less relevant labels (e.g., UI types) and uniformly map the remaining labels to a common label space containing 13 classes: ‘Checkbox’, ‘Button’, ‘Container’, ‘Dialog’, ‘Icon’, ‘PageControl’, ‘Picture’, ‘SegmentedControl’, ‘Slider’, ‘TabBar’, ‘Text’, ‘TextField’, and ‘Toggle’, and obtain a multi-platform uniformed dataset with raw UI widget annotations. We provide the original label statistics and their converted labels in Appendix F.
We name the above collected screenshot dataset the Core-set, which we will use to construct the elementary and advanced task data. Besides, we also employ third-party training datasets to enrich our data source and avoid overfitting our predefined tasks. A complete statistics of the training dataset of Ferret-UI 2 is summarized in Table 1, indicating that the dataset distribution is very unbalanced across different platforms. In particular, the number of iPad and AppleTV screenshots is significantly smaller than that of other platforms. To address this, we ($i$) assign different loss weights to different platforms during training, and ($ii$) generate all three types of advanced tasks for each example of iPad and AppleTV platforms and generate only 1 type of advanced task for each example of other platforms.
@@
For advanced tasks, we prompt GPT-4o with bounding box annotations of a given screenshot and require GPT-4o to generate QA tasks related to the UI widgets in the screenshot. Unlike Ferret-UI, which mainly focuses on spatial descriptions due to the limitation of using textual prompts without image information (i.e., screenshots) for bounding box annotations, Ferret-UI 2 leverages GPT-4o to generate advanced task data that covers a variety of aspects of UI understanding. This is possible because GPT-4o demonstrates an improved ability to comprehend the spatial relationships between UI widgets when provided with the screenshot as input. Specifically, we prompt GPT-4o to generate 3 types of advanced tasks (shown in Figure 2) including: ($i$) comprehensive description: describe global and local functionalities of the screen, ($ii$) multi-round perception QA: multi-round question answering regarding the UI perception capabilities, and ($iii$) multi-round interaction QA: multi-round question answering regarding the single-step and user-centric UI interactions based on the current screen status. More detailed requirements and prompts for GPT-4o when generating advanced tasks are provided in Appendix B.
We empirically find it hard for GPT-4o to find the location of referred UI widgets with the original screenshot as input (i.e., bad grounding capability). To address this, we use Set-of-Mark (SoM) visual prompting (Yang et al., 2023) when generating multi-round perception and interaction QA training samples. One example of SoM prompting and its generated data sample is shown in Figure 3, where each UI widget is marked with a corner-style bounding box and a unique number tag for easy identification. Furthermore, the same class of UI widgets have the same color for visual prompting to help GPT-4o better differentiate the bounding boxes of spatially close or nested widgets. Please refer to Figure 5 for visual prompts on other platforms, and Appendix C for additional examples of generated data for advanced tasks.
@@ 3.2 Model Architecture
Adaptive Gridding. Local image features are extracted by calculating the optimal grid size using our proposed adaptive $N$-gridding mechanism, then resizing and encoding the visual features of each grid. This is a key model innovation compared to Ferret-UI. As shown in Algorithm 1, the optimal gridding size $N_{w}$ and $N_{h}$ is determined when the gridding and resizing based on $(N_{w},N_{h})$ lead to minimal aspect ratio change times the relative pixel number change, under the constraint $N_{w}+N_{h}\leq N$, where $N$ is the size limit. With the size limit $N$, the total number of grids is upper bounded by $\lfloor\frac{N^{2}}{4}\rfloor$. Compared to the AnyRes module that has unbounded cost, the key differences of adaptive $N$-gridding is it automatically finds the optimal gridding configuration, i.e., least resolution distortion (aspect ratio change and pixel number change) within a predefined inference cost limit that $N_{w}\times N_{h}\leq\lfloor\frac{N^{2}}{4}\rfloor$, which is both information-preserving and efficient for local encoding. With adaptive gridding, Ferret-UI 2 understands UI screens and provides user-centered interactions with an optimal configuration at any resolution given the inference cost limit specified as size limit $N$.
@@ 4.1 Experiment Setup
Table 2: Results on our constructed benchmarks for elementary and advanced tasks, as well as the GUIDE benchmark (Chawla et al., 2024). Results on elementary and advanced tasks are averaged over all platforms, including iPhone, Android, iPad, Webpage, and AppleTV. Each platform includes 6 elementary tasks and 3 advanced tasks. ($\dagger$) In tasks that require referring, GPT-4o is equipped with set-of-mark (SoM) prompting by adding a red rectangular box to screenshots for the referred widget. Note that SoM visual prompting is not used for Ferret-UI and Ferret-UI 2.
@@
Model. Following Ferret-UI (You et al., 2024), Ferret-UI 2 uses a CLIP ViT-L/14 model as the image encoder; for the LLM backone, besides Vicuna-13B (Chiang et al., 2023) as used in the original Ferret-UI, we also tried 2 additional LLMs at mobile scales, including Gemma-2B (Team et al., 2024) and Llama3-8B (Dubey et al., 2024). As to dynamic high-resolution image encoding, we set the size limit $N$ to 8, so that the maximal grid number is 16 for adaptive gridding.
Evaluation. At a high level, model evaluation falls into two broad categories: ($i$) benchmarks we constructed, and ($ii$) public benchmarks. For our benchmarks, we created a total of 45, including 6 elementary tasks and 3 advanced tasks per platform type, across 5 platforms. For elementary tasks, we follow the evaluation metrics outlined by You et al. (2023). For advanced tasks, we use GPT-4o to score generated answers for a given screenshot and user query, visually prompting GPT-4o with a red rectangular bounding box. Advanced tasks are tested using GPT-4o evaluation score and multi-IoU. The multi-IoU is calculated by first matching predicted bounding boxes with ground truth bounding boxes and then calculate the average IoU of each pair of bounding boxes ($\text{IoU}=0$ if no match). Furthermore, we conduct next-action prediction test given previous action history on the GUIDE benchmark (Chawla et al., 2024), and evaluate the semantic similarity w.r.t. the reference answer and grounding Intersection-over-Union (IoU). Additionally, we evaluate our model on the recently released GUI-World benchmark (Chen et al., 2024a) on the supported platforms following the original GPT-4 evaluation protocol as in Chen et al. (2024a), which does not include evaluating the grounding capability for interaction-related UI tasks.
@@ 4.2 Experiment Results
  • Ferret-UI 2, powered by Llama-3-8B, delivers the best results across most metrics. It achieves the highest GPT-4o score on advanced tasks, with a notable 89.73, surpassing Ferret-UI by 43.92 points and GPT-4o by 12.0 points. Notably, Ferret-UI 2 with Llama-3-8B also achieves the highest IoU score on the GUIDE benchmark with 55.78, indicating superior grounding capability over the other models.
@@
Results on GUI-World. To further demonstrate the zero-shot performance of using Ferret-UI 2 out of the box, we further test our model on the recently released GUI-World benchmark (Chen et al., 2024a). Results are summarized in Table 3. Clearly, Ferret-UI 2 does not overfit to the training data, and can generalize well to the test data in the wild. Notably, Ferret-UI 2 outperforms GUI-Vid (Chen et al., 2024a), a model developed in the GUI-World paper, on supported platforms including iOS, Android and Webpages.
@@ 4.3 Ablation Study
Cross-Platform Transferability. In Table 4, we evaluate the zero-shot platform (domain) transfer capability of the Ferret-UI 2 model by training it on elementary tasks from one platform and testing it on other platforms. These results provide insights into how well the model generalizes across different platforms. We observe similar performance patterns across both referring and grounding tasks. Specifically,
  • iPhone transfers well to iPad and Android platforms on both tasks, achieving at least 68.1 and 71.2 scores on referring tasks and 65.2 and 63.1 scores on grounding tasks due to its diverse screenshot contents (because of the large screenshot number) and similar resolutions and aspect ratios with other two platforms. iPad and Android can also transfer fairly well to the iPhone domain, all achieving around 65 scores.
  • AppleTV and Web do not transfer very well to the mobile domains, including iPhone, Android, and iPad, achieving the highest referring score of 59.2 and grounding score of 54.0, possibly because they are mostly landscape screenshots, which is in contrast to the mostly portrait screenshots in mobile platforms. Models trained on other domains all achieve poor performance on the AppleTV test domain, with the highest score of around 40 on both kinds of tasks, which is reasonable since there is a large domain gap in terms of AppleTV’s content distribution compared to other domains.
The results suggest that ($i$) iPhone, iPad and Android have similar content distribution, which help them generalize to each other; ($ii$) models trained on more diverse contents (e.g., around 100k iPhone data) can generalize better to other platforms; ($iii$) platforms with similar resolution and aspect ratios may transfer better to each other; and ($iv$) good transferability among some of these platforms contributes to the cross-platform performance of Ferret-UI 2.
@@
Training Data Model Test Data
iPhone v1 iPhone v2
GPT4o Score Multi-IoU GPT4o Score Multi-IoU
@@
Ablation on Architecture and Dataset Improvements. Table 5 presents a comparison between the Ferret-UI and Ferret-UI 2 model trained on different versions of the iPhone dataset. Models are evaluated on the test set corresponding to advanced tasks. The results indicate that both the architectural enhancements, particularly the adaptive $N$-gridding, and the improved dataset (iPhone v2) contribute to performance gains.
Specifically, when evaluated on the iPhone v1 test set, Ferret-UI 2 shows a slight improvement over Ferret-UI, with the GPT4o score increased from 91.3 to 93.7 and the Multi-IoU score increased from 36.89 to 37.12. However, the improvements are more pronounced on the iPhone v2 dataset. Here, Ferret-UI 2 achieves a GPT4o score of 89.7, outperforming Ferret-UI’s 85.97, along with a substantial Multi-IoU score boost from 39.81 to 41.73. These results suggest that while both architecture and dataset enhancements contribute to overall performance, the new dataset plays a more significant role in driving improvements, particularly on more challenging tasks.

### v2

@@ 1 Introduction
One prominent effort in this space is Ferret-UI (You et al., 2024), which has advanced the field of referring and grounding UIs. However, though taking an any-resolution approach (Liu et al., 2024a), Ferret-UI is constrained by a fixed grounding resolution (i.e., 336$\times$672 and 672$\times$336), and focuses on single-type platforms (i.e., mobile devices including iPhone and Android), limiting its applicability in the context of today’s highly diverse platform landscape. For example, as illustrated in Figure 1, one notable difference among these four exemplified platforms is resolution, the native resolution of iPhone differs from that of iPad, Web UI, and also AppleTV, directly applying Ferret-UI across these platforms presents significant challenges. Another major challenge is the lack of platform-specific, high-quality data, given different platforms. Though Ferret-UI’s approach to training data generation can be extended to these platforms, it primarily relies on text-based GPT-4 prompting, where bounding boxes are represented in a purely textual format. This absence of visual input and spatial relationships between UI elements diminishes the quality of training data, which in turn limits the performance and effectiveness of the resulting model.
Figure 1: Real examples of a single Ferret-UI 2 model interacting with four different platforms (iPhone, iPad, Webpage, and AppleTV) for UI understanding. Refer to Appendix H for more examples including multi-step interactions.
@@
Second, Ferret-UI 2 supports high-resolution image encoding via the any-resolution method (Liu et al., 2024a; Zhang et al., 2024c). However, going beyond that, we introduce an enhanced adaptive gridding approach to maintain perception capabilities at the original resolution of the UI screenshot, ensuring more accurate recognition of visual elements. By leveraging human-collected bounding box annotations, we enhance referring and grounding precision, which allows for a more detailed understanding of UI components and their relationships.
@@ 2 Related Work
Multi-Platform UI Agents. Multi-platform UI agents have emerged to address the growing complexity of device ecosystems, supporting a variety of devices and platforms, including mobile, web, and desktop environments. Zheng et al. (2024a); Cheng et al. (2024) demonstrated GPT-4V as a generalist agent when grounded. Recent works like CogAgent (Hong et al., 2023) support UI navigation on both PC webpages and Android devices. Mobile-Agent V2 (Wang et al., 2024a) features Harmony OS and Android OS for non-English and English scenarios. Ferret-UI (You et al., 2024) focuses on mobile UI understanding for both Android and iPhone screenshots using multimodal LLMs (McKinzie et al., 2024; Zhang et al., 2024b), with a focus on referring and grounding capabilities. These agents aim to perform more complex, user-intent-based interactions across multiple device types, paving the way for truly generalist multimodal agents.
UI-Agent Benchmarks. The evaluation of UI agents requires specialized benchmarks to test various aspects of UI interaction, including task execution, navigation, and interaction understanding. Rico (Deka et al., 2017) remains a foundational dataset for mobile app interaction, while Mobile-Env (Zhang et al., 2023), AndroidEnv (Toyama et al., 2021), AndroidWorld (Rawles et al., 2024a), Android in-the-Wild (Rawles et al., 2024b), AndroidControl (Li et al., 2024b), and AMEX (Chai et al., 2024) provide benchmarks for mobile device control. Windows Agent Arena (Bonatti et al., 2024) introduces a benchmark focusing on PC Windows environment. OSWorld (Xie et al., 2024) takes a broader approach by providing benchmarks for agents in real computer environments, including Ubuntu, MacOS, and Windows. Web-based interaction benchmarks include WebSRC (Chen et al., 2021), Mind2Web (Deng et al., 2024), and WebCanvas (Pan et al., 2024) focusing on structural reading comprehension and task execution in web environments. OmniACT (Kapoor et al., 2024) supports both desktop and web interfaces. More recently, MobileAgentBench (Wang et al., 2024c) and VisualWebBench (Liu et al., 2024b) have introduced taxonomies designed to evaluate the performance of multimodal agents across both mobile and web interfaces. VisualAgentBench (Liu et al., 2024c) expands this with a focus on multimodal LLMs as visual foundation agents. GUI Odyssey (Lu et al., 2024) provides benchmarks for cross-app navigation. GUI-World (Chen et al., 2024a) pioneer in covering multi-platform benchmarking, while CRAB (Xu et al., 2024) further tests cross-environment tasks for GUI agents. These benchmarks contribute to a growing need for unified, multi-platform evaluations that can assess UI agents’ adaptability, precision, and efficiency.
@@ 3.1 Dataset Construction
  • iPhone, iPad, and AppleTV: We use human-collected iPhone, iPad, and AppleTV data under diverse usage scenarios and human-annotated widget bounding box coordinates and labels. To save annotation costs, we do not collect text annotations; instead, text bounding boxes are replaced by screen-wide OCR-detected text and bounding boxes using an OCR confidence threshold of 0.5.
@@
  • Android: The Android data for screenshots, bounding boxes, and text annotations is transformed from the RICO dataset (Deka et al., 2017). Similar to the WebUI dataset, we also perform picture-only OCR to complete the missing text annotations in picture widgets.
Use of this data for the purpose of this paper was approved by our organization’s legal team.
@@
Table 1: A summary of datasets across various platforms used to train Ferret-UI 2. The screenshot resolution statistics are shown in Appendix F.
@@
Figure 3: Example of set-of-mark visual prompting (Yang et al., 2023) (left) and one of its generated advanced task training examples (right).
Despite the different types of label spaces from various sources, we filter out bounding boxes associated with less relevant labels (e.g., UI types) and uniformly map the remaining labels to a common label space containing 13 classes: ‘Checkbox’, ‘Button’, ‘Container’, ‘Dialog’, ‘Icon’, ‘PageControl’, ‘Picture’, ‘SegmentedControl’, ‘Slider’, ‘TabBar’, ‘Text’, ‘TextField’, and ‘Toggle’, and obtain a multi-platform uniformed dataset with raw UI widget annotations. We provide the original label statistics and their converted labels in Appendix G.
We name the above-collected screenshot dataset the Core-set, which we will use to construct the elementary and advanced task data. Besides, we also employ third-party training datasets to enrich our data source and avoid overfitting our predefined tasks. A complete statistics of the training dataset of Ferret-UI 2 is summarized in Table 1, indicating that the dataset distribution is very unbalanced across different platforms. In particular, the number of iPad and AppleTV screenshots is significantly smaller than that of other platforms. To address this, we ($i$) assign different loss weights to different platforms during training, and ($ii$) generate all three types of advanced tasks for each example of iPad and AppleTV platforms and generate only 1 type of advanced task for each example of other platforms.
@@
For advanced tasks, we prompt GPT-4o with bounding box annotations of a given screenshot and require GPT-4o to generate QA tasks related to the UI widgets in the screenshot. Unlike Ferret-UI, which mainly focuses on spatial descriptions due to the limitation of using textual prompts without image information (i.e., screenshots) for bounding box annotations, Ferret-UI 2 leverages GPT-4o to generate advanced task data that covers a variety of aspects of UI understanding. This is possible because GPT-4o demonstrates an improved ability to comprehend the spatial relationships between UI widgets when provided with the screenshot as input. Specifically, we prompt GPT-4o to generate 3 types of advanced tasks (shown in Figure 2) including: ($i$) comprehensive description: describe global and local functionalities of the screen, ($ii$) multi-round perception QA: multi-round question answering regarding the UI perception capabilities, and ($iii$) multi-round interaction QA: multi-round question answering regarding the single-step and user-centric UI interactions based on the current screen status. More detailed requirements and prompts for GPT-4o when generating advanced tasks are provided in Appendix C.
We empirically find it hard for GPT-4o to find the location of referred UI widgets with the original screenshot as input (i.e., bad grounding capability). To address this, we use Set-of-Mark (SoM) visual prompting (Yang et al., 2023) when generating multi-round perception and interaction QA training samples. One example of SoM prompting and its generated data sample is shown in Figure 3, where each UI widget is marked with a corner-style bounding box and a unique number tag for easy identification. Furthermore, the same class of UI widgets have the same color for visual prompting to help GPT-4o better differentiate the bounding boxes of spatially close or nested widgets. Please refer to Figure 5 for visual prompts on other platforms, and Appendix D for additional examples of generated data for advanced tasks.
@@ 3.2 Model Architecture
Adaptive Gridding. Local image features are extracted by calculating the optimal grid size using our proposed adaptive $N$-gridding mechanism, then resizing and encoding the visual features of each grid. This is a key model innovation compared to Ferret-UI. As shown in Algorithm 1, the optimal gridding size $N_{w}$ and $N_{h}$ is determined when the gridding and resizing based on $(N_{w},N_{h})$ lead to minimal aspect ratio change times the relative pixel number change, under the constraint $N_{w}+N_{h}\leq N$, where $N$ is the size limit. With the size limit $N$, the total number of grids is upper bounded by $\lfloor\frac{N^{2}}{4}\rfloor$. Compared to the AnyRes module that has unbounded cost, the key differences of adaptive $N$-gridding is it automatically finds the optimal gridding configuration, i.e., least resolution distortion (aspect ratio change and pixel number change) within a predefined inference cost limit that $N_{w}\times N_{h}\leq\lfloor\frac{N^{2}}{4}\rfloor$, which is both information-preserving and efficient for local encoding. With adaptive gridding, Ferret-UI 2 understands UI screens and provides user-centered interactions with an optimal configuration at any resolution given the inference cost limit specified as $N$.
@@ 4.1 Experiment Setup
Table 2: Results on our constructed benchmarks for elementary and advanced tasks, as well as the GUIDE benchmark (Chawla et al., 2024). Results on elementary and advanced tasks are averaged over all platforms, including iPhone, Android, iPad, Webpage, and AppleTV. Each platform includes 6 elementary tasks and 3 advanced tasks. SeeClick model (Cheng et al., 2024) trained on their original data is compared. ($\dagger$) In tasks that require referring, GPT-4o is equipped with set-of-mark (SoM) prompting by adding a red rectangular box to screenshots for the referred widget. Note that SoM visual prompting is not used for Ferret-UI and Ferret-UI 2.
@@
Model. Following Ferret-UI (You et al., 2024), Ferret-UI 2 uses a CLIP ViT-L/14 model as the image encoder; for the LLM backbone, besides Vicuna-13B (Chiang et al., 2023) as used in the original Ferret-UI, we also tried 2 additional LLMs at mobile scales, including Gemma-2B (Team et al., 2024) and Llama3-8B (Dubey et al., 2024). As to dynamic high-resolution image encoding, we set the size limit $N$ to 8, so that the maximal grid number is 16 for adaptive gridding.
Evaluation. At a high level, model evaluation falls into two broad categories: ($i$) benchmarks we constructed, and ($ii$) public benchmarks. For our benchmarks, we created a total of 45, including 6 elementary tasks and 3 advanced tasks per platform type, across 5 platforms. For elementary tasks, we follow the evaluation metrics outlined by You et al. (2023). For advanced tasks, we use GPT-4o to score generated answers for a given screenshot and user query, visually prompting GPT-4o with a red rectangular bounding box. Advanced tasks are tested using GPT-4o evaluation score and multi-IoU. The multi-IoU is calculated by first matching predicted bounding boxes with ground truth bounding boxes and then calculating the average IoU of each pair of bounding boxes ($\text{IoU}=0$ if no match). Furthermore, we conduct a next-action prediction test given previous action history on the GUIDE benchmark (Chawla et al., 2024), and evaluate the semantic similarity w.r.t. the reference answer and grounding Intersection-over-Union (IoU). Additionally, we evaluate our model on the recently released GUI-World benchmark (Chen et al., 2024a) on the supported platforms following the original GPT-4 evaluation protocol as in Chen et al. (2024a), which does not include evaluating the grounding capability for interaction-related UI tasks.
@@ 4.2 Experiment Results
  • Ferret-UI 2, powered by Llama-3-8B, delivers the best results across most metrics. It achieves the highest GPT-4o score on advanced tasks, with a notable 89.73, surpassing Ferret-UI by 43.92 points and GPT-4o by 12.0 points. Notably, Ferret-UI 2 with Llama-3-8B also achieves the highest IoU score on the GUIDE benchmark with 55.78, indicating superior grounding capability.
@@
Results on GUI-World. To further demonstrate the zero-shot performance of using Ferret-UI 2 out of the box, we further test our model on the recently released GUI-World benchmark (Chen et al., 2024a). Results are summarized in Table 3. Clearly, Ferret-UI 2 does not overfit the training data and can generalize well to the test data in the wild. Notably, Ferret-UI 2 outperforms GUI-Vid (Chen et al., 2024a), a model developed in the GUI-World paper, on supported platforms including iOS, Android, and Webpages.
@@ 4.3 Ablation Study
Cross-Platform Transferability. In Table 4, we evaluate the zero-shot platform (domain) transfer capability of the Ferret-UI 2 model by training it on elementary tasks from one platform and testing it on other platforms. These results provide insights into how well the model generalizes across different platforms. We observe similar performance patterns across tasks. Specifically,
  • iPhone transfers well to iPad and Android platforms on both tasks, achieving at least 68.1 and 71.2 scores on referring tasks and 65.2 and 63.1 scores on grounding tasks due to its diverse screenshot contents (because of the large screenshot number) and similar resolutions and aspect ratios with other two platforms. iPad and Android also transfer fairly well to iPhone at around 65 scores.
  • AppleTV and Web do not transfer very well to the mobile domains, including iPhone, Android, and iPad, achieving the highest referring score of 59.2 and grounding score of 54.0, possibly because they are mostly landscape screenshots, which is in contrast to the mostly portrait screenshots in mobile platforms. Models trained on other domains all achieve poor performance on the AppleTV test domain, with the highest score of around 40 on both kinds of tasks, which is reasonable due to a large domain gap in terms of AppleTV’s content distribution compared to other domains.
The results suggest that ($i$) iPhone, iPad and Android have similar content distribution, which helps them generalize to each other; ($ii$) models trained on more diverse contents (e.g., around 100k iPhone data) can generalize better to other platforms; ($iii$) platforms with similar resolution and aspect ratios may transfer better to each other; and ($iv$) good transferability among some of these platforms contributes to the cross-platform performance of Ferret-UI 2.
@@
Training Data Model iPhone v1 iPhone v2
GPT-4o Score Multi-IoU GPT-4o Score Multi-IoU
@@
Ablation on Architecture and Dataset Improvements. Table 5 presents a comparison between the Ferret-UI and Ferret-UI 2 model trained on different versions of the iPhone dataset. Models are evaluated on the test set corresponding to advanced tasks. The results indicate that both the adaptive $N$-gridding and improved dataset (iPhone v2) contribute to performance gains.
Specifically, when evaluated on the iPhone v1 test set, Ferret-UI 2 shows a slight improvement over Ferret-UI, with the GPT-4o score increased from 91.3 to 93.7 and the Multi-IoU score increased from 36.89 to 37.12. However, the improvements are more pronounced on the iPhone v2 dataset. Here, Ferret-UI 2 achieves a GPT-4o score of 89.7, outperforming Ferret-UI’s 85.97, along with a substantial Multi-IoU score boost from 39.81 to 41.73. These results suggest that while both architecture and dataset enhancements contribute to overall performance, the new dataset plays a more significant role in driving improvements, particularly on more challenging tasks.